40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 22 23:06:13 2011, kiwamu, Update, SUS, MC watchdogs 
    Reply  Fri Nov 25 20:41:48 2011, rana, Update, SUS, MC watchdogs 
       Reply  Thu Dec 1 17:33:18 2011, kiwamu, Update, SUS, watchdogs fixed 
Message ID: 6060     Entry time: Thu Dec 1 17:33:18 2011     In reply to: 6010
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: watchdogs fixed 

The watchdogs' issue has been solved and they are now working fine.

It was just because one of the Sorensens had been off.

The Sorensen is the one supplying +5 V in the 1X5 rack.
This +5 V is actually used as a pull-up-current to properly drive the MAX333As (CMOS analog switch) in the coil drivers (D010001).
So this was it.

Quote from #6010

Tonight we noticed that, in fact, the watchdogs don't work for any of the corner optics (I confirmed that they do work for the ETMs).

ELOG V3.1.3-